Please help me answer the following:
(A) Find the binomial probability P(x = 4), where n = 12 and p = 0.40.
(B) Set up, without solving, the binomial probability P(x is at most 4) using probability notation.
(C) How would you find the normal approximation to the binomial probability P(x = 4) in part A? Please show how you would calculate µ and σ in the formula for the normal approximation to the binomial, and show the final formula you would use without going through all the calculations.

Answer by stanbon(57307)

(A) Find the binomial probability P(x = 4), where n = 12 and p = 0.40.
P(x=4) = 12C4*0.4^4*0.6^8 = 0.2128
-------------------------------------------------
(B) Set up, without solving, the binomial probability P(x is at most 4) using probability notation.
P(0<= x <=4) = 12C0*0.4^0*0.6^12 + 12C1*0.4*0.6^11+...+12C4*0.4^4*0.6^8
===============================================================================
(C) How would you find the normal approximation to the binomial probability P(x = 4) in part A?
Please show how you would calculate µ and σ in the formula for the normal approximation to the binomial, and show the final formula you would use without going through all the calculations.
u = np = 0.4*12 = 4.8
s = sqrt(npq) = sqrt(4.8*0.6) = 1.6971
-----
Normal approx: P(3.5< x < 4.5)
z(3.5) = (3.5-4.8)/1.6971
z(4.5) = (4.5-4.8)/1.6871
---
P(z(3.5)<= z =================================================
cheers,
Stan H.
=================